# RoboArmGS: High-Quality Robotic Arm Splatting via Bézier Curve Refinement

### v1

@@ 3.3 Bézier-based Motion Refiner
While existing methods directly apply Forward Kinematics (FK) transformations without considering the motion process, we introduce a Bézier-based Motion Refiner (BMR) to correct for inevitable discrepancies between idealized FK models and real-world dynamics. It employs a hierarchical decomposition, combining a learnable, time-varying global correction with static, per-joint offsets. This design enforces temporal smoothness globally while allowing for local kinematic flexibility.
@@
The time-varying global correction, $\mathbf{T}_{\text{Bézier}}(t)$, is modeled by a learnable Bézier curve that operates in a 9-dimensional parameter space. We parameterize the residual transformations as 9D vectors $\boldsymbol{\delta}=[\Delta\mathbf{x},\Delta\mathbf{r}]$, consisting of a 3D translation $\Delta\mathbf{x}$ and a continuous 6D rotation representation $\Delta\mathbf{r}$ [zhou2019continuity]. The Bézier curve, defined by $K+1$ learnable control points $\{\mathbf{p}_{k}\}_{k=0}^{K}\subset\mathbb{R}^{9}$, outputs a 9D residual vector for any time $t$:
@@
where $B_{k}^{K}(t)$ are the Bernstein basis polynomials. We use a high-order curve ($K=19$) to capture complex motion patterns and scale the output by an influence factor $\omega$ to ensure it acts as a subtle refinement. The resulting 9D vector is then converted to a valid SE(3) matrix $\mathbf{T}_{\text{Bézier}}(t)$ via Gram-Schmidt orthogonalization for the rotational part.
@@
The static, per-joint offset, $\mathbf{T}_{\text{embed}}^{(j)}$, is represented by a learnable embedding $\mathbf{e}^{(j)}\in\mathbb{R}^{9}$ for each moving joint $j$. These time-invariant embeddings, also parameterized as 9D vectors and initialized to zero, are optimized to correct for local, static errors such as link length discrepancies or joint zero-position offsets. By operating at the joint level, our model directly targets the primary sources of kinematic error within the robot’s articulated structure.
@@
The final refined pose for each joint is computed by sequentially composing the baseline FK pose from MuJoCo with our two learned corrections. The composition order is crucial: the FK pose is first modified by the learnable Bézier correction, and then further refined by the static per-joint offset corresponding to the joint that actuates the joint’s link:
$\mathbf{T}_{\text{final}}^{(j)}(t)=\mathbf{T}_{\text{FK}}^{(g)}(t)\circ\mathbf{T}_{\text{Bézier}}(t)\circ\mathbf{T}_{\text{embed}}^{(j)}$ (7)
@@ 3.4 Optimization and Regularization
The primary supervision signal is a rendering loss comparing the rendered images with ground truth frames. Following standard practice [kerbl20233d], we use a combination of an L1 photometric loss and a D-SSIM term:
@@
To ensure the optimized Gaussians remain geometrically plausible and well-aligned with their parent faces, we adopt two regularization terms inspired by GaussianAvatars [qian2024gaussianavatars]. We penalize the local position offset $\mu_{j}$ and scale $s_{j}$ if they exceed the thresholds, preventing primitive drifting and visual jittering during animation:
@@ 4 Experiments
[…]
Table 1: Quantitative comparison with state-of-the-art methods. Our model demonstrates superior performance in both static Novel-View Synthesis (left) and dynamic Novel-Pose Synthesis (right). Green indicates the best and yellow indicates the second-best performance.
@@ 4.1 Setup
Existing public datasets for robotics typically provide either visual, linguistic, or kinematic data in isolation, making them insufficient for the task of building high-fidelity dynamic digital assets. To address this gap, we introduce RoboArm4D, the first public benchmark specifically designed for creating and evaluating dynamic digital assets of robotic arms. Our dataset features several widely-used industrial arms, such as the Franka Research 3, UR-5e, and ABB IRB 120. For each arm, we offer a complete, synchronized package that includes high-resolution monocular videos, calibrated camera parameters, time-stamped joint trajectories, and the associated URDF file. Additional information regarding the data collection process and the package contents can be found in the supplementary material.
@@
Our model, RoboArmGS, is implemented in PyTorch and trained for 600,000 iterations on a single NVIDIA H100 GPU using the Adam optimizer. For 3D Gaussian attributes, we follow the standard learning rate schedule from 3DGS [kerbl20233d], with position learning rates exponentially decayed to 1% of initial values. The BMR parameters are optimized with a learning rate of 0.0015, weight decay of 0.0001, and influence factor $\omega=0.1$. Gaussians are initialized by uniform sampling on the kinematic mesh surface, with binding-aware adaptive densification activated every 100 iterations from iteration 500 to 60,000; unstable Gaussian opacities are reset every 3,000 iterations to prune floaters. FK is computed via MuJoCo [todorov2012mujoco] engine, and regularization hyperparameters are set as $\lambda_{\text{pos}}=0.01,\epsilon_{\text{pos}}=1.0$, $\lambda_{\text{scale}}=1.0,\epsilon_{\text{scale}}=0.6$, and $\lambda_{\text{vel}}=0.001$.
Figure 3: Qualitative comparison on novel-view synthesis. Our RoboArmGS achieves significantly sharper and more photorealistic renderings than Robo-GS [lou2025robo], with better geometric accuracy and texture preservation.
Figure 4: Qualitative comparison on novel-pose synthesis. RoboArmGS achieves photorealistic rendering with precise geometric alignment. Baselines fail in different ways: 4DGS and Deformable 3DGS produce blurred, distorted results lacking kinematic control; 3DGS+FK, despite correct Gaussian positioning, generates severe artifacts from inaccurate motion. Our BMR module successfully addresses both motion accuracy and rendering coherence.
@@ 4.2 Main Results
We first evaluate the static 3D reconstruction quality on the novel-view synthesis task. As shown in Tab. 1, RoboArmGS demonstrates superior performance, achieving the best results across all metrics, which confirms the high fidelity and perceptual realism of our renderings. We attribute this success to our SGB module, which effectively leverages the robot’s kinematic mesh as a powerful geometric prior. By constraining the Gaussians to this structure, our model produces a more coherent 3D representation, establishing a high-quality foundation for dynamic modeling. Fig. 3 provides a qualitative comparison with Robo-GS [lou2025robo], clearly demonstrating the superior visual quality of our approach.
@@
We evaluate our dynamic modeling capability on the novel-pose synthesis task, which renders the robot in arbitrary, unseen joint configurations. As shown in Tab. 1, RoboArmGS substantially outperforms all baselines across all metrics. Critically, dynamic scene baselines model [wu20244d, yang2024deformable] temporal evolution using temporal indices rather than kinematic parameters, thus lacking the ability to render from specified joint configurations which is essential for digital twin applications. Even on held-out test frames, their generic deformation models fail to capture articulated motions, producing significant artifacts. The 3DGS + FK baseline, despite correct kinematic positioning, suffers from severe rendering artifacts and structural distortions. Our success stems from the Bézier-based Motion Refiner, which bridges the kinematic modeling gap through temporally smooth global corrections and per-joint offsets, ensuring kinematic accuracy and rendering coherence. Fig. 4 demonstrates our model renders sharp, structurally coherent images precisely aligned with ground truth, validating RoboArmGS as a controllable, motion-accurate digital asset.
@@ 4.3 Ablation Study
w/o SGB 20.452 0.809 0.438 w/o BMR 23.484 0.922 0.075
    - w/o Adaptive Densification 20.414 0.803 0.451     - w/o Bézier Correction 22.161 0.903 0.089
@@
We validate the effectiveness of each proposed component through comprehensive ablation studies, presented in Tab. 2. For Novel-View Synthesis, removing the SGB module severely degrades performance. Furthermore, ablating its sub-components confirms their necessity for high-fidelity static reconstruction. For Novel-Pose Synthesis, removing the BMR module demonstrates its critical role in bridging the kinematic modeling gap. As illustrated in Fig. 6, the impact of BMR is motion-dependent: the base region (green box), which undergoes minimal motion, shows negligible differences between the full model and the ablated variant, while the gripper region (red box), which experiences large-scale motion, exhibits significant visual artifacts without BMR. Further analysis reveals that both the Learnable Bézier Correction for global dynamic errors and the Per-Joint Static Offset for local static errors are essential and complementary. This study empirically validates that all proposed components of SGB and BMR are indispensable to the success of RoboArmGS.
Figure 6: Ablation study visualization. Green box: base region with minimal motion shows similar results for both variants. Red box: gripper region with significant motion reveals rendering artifacts when BMR is removed, demonstrating its importance for high-mobility components.
@@
We investigate the impact of the influence factor $\omega$, which scales our learnable Bézier correction, with results presented in Tab. 3. Setting $\omega=0.0$ is equivalent to removing the global correction, which results in the lowest performance and confirms the necessity of this component. As we introduce a small correction factor ($\omega=0.01$), performance dramatically improves across all metrics. Our chosen value of $\omega=0.1$ achieves the best PSNR and SSIM scores, striking an optimal balance for overall motion fidelity. However, further increasing $\omega$ to 1.0 leads to a slight degradation in performance, suggesting that an overly strong correction can introduce instability or overfitting to the training trajectory. This study validates our selection of $\omega=0.1$ as an effective and stable choice for motion refinement.
@@ 5 Conclusion
We introduce RoboArmGS, a hybrid representation that couples Structured Gaussian Binding with a Bézier-based Motion Refiner to jointly ensure geometric consistency and accurate real-world motion modeling. This design effectively resolves long-standing limitations, including rendering artifacts and URDF–rigged motion mismatch in the robot arm reconstruction. We further release RoboArm4D, the first benchmark dedicated to dynamic robotic arm digital asset reconstruction. Comprehensive experiments show that RoboArmGS achieves state-of-the-art novel-view and novel-pose rendering fidelity.
[…]
Supplementary Material
@@ B.2.1 Camera Pose Calibration
Accurately determining the pose of a static camera is crucial. To achieve this, we first captured a short calibration video (approx. 30 seconds) where the handheld camera was moved extensively around the static robotic arm and its environment. This multi-view sequence was processed using VGGT [wang2025vggt] to generate a sparse 3D reconstruction of the scene and to precisely compute the camera poses for each frame of the calibration video. From this set of calibrated poses, we selected a single, fixed viewpoint for the subsequent motion capture. This process effectively pre-calibrates the static camera’s extrinsic parameters within the scene’s coordinate frame.
@@ B.2.3 Data Processing Pipeline
  • Camera Pose Estimation: We used VGGT [wang2025vggt] to estimate the camera intrinsics and the per-frame extrinsic poses (camera-to-world transformation). Poses were optimized over the entire sequence to ensure global consistency.
  • Foreground Segmentation: To separate the robotic arm from the background, we employed the Segment Anything Model 2 (SAM2) [ravi2024sam]. We provided a few initial keyframe masks, and SAM2 [ravi2024sam] automatically propagated the segmentation to the entire sequence, followed by minor manual refinement where necessary. This resulted in a pixel-perfect foreground mask for each frame.
@@ Appendix C Additional Experiments Results
In this section, we provide a detailed per-scene quantitative analysis of RoboArmGS on the RoboArm4D dataset, specifically focusing on the Universal Robots UR5e and ABB IRB 120 sequences. As shown in Tab. 1, our method achieves consistently high performance across different robotic morphologies, validating the generalizability of our proposed framework.
Table 1: Per-scene Quantitative Results on RoboArm4D. Detailed performance of RoboArmGS on Universal Robots UR5e and ABB IRB 120. The results are reported separately for Novel-View Synthesis and Novel-Pose Synthesis tasks.
Robot PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Novel-View Synthesis
Universal Robots UR5e 34.92 0.968 0.026
ABB IRB 120 35.02 0.969 0.038
Novel-Pose Synthesis
Universal Robots UR5e 36.25 0.984 0.016
ABB IRB 120 37.36 0.982 0.018

### v2

@@ 3.3 Bézier-based Motion Refiner
To reconcile idealized kinematic models with physical reality, we introduce the Bézier-based Motion Refiner (BMR). While standard FK provides a strong motion prior, they often fail to account for real-world deviations such as control latency or joint friction. Drawing inspiration from the smooth temporal parameterization in BézierGS (Ma et al., 2025), our BMR models these discrepancies as a continuous residual field. We employ a hierarchical decomposition that couples a learnable, time-varying global correction with static, per-joint offsets. This design leverages the inherent smoothness of Bézier curves to ensure global temporal coherence while providing the local flexibility necessary to compensate for systematic kinematic errors.
@@
The time-varying global correction, $\mathbf{T}_{\text{Bézier}}(t)$, is modeled by a learnable Bézier curve that operates in a 9-dimensional parameter space. Unlike BézierGS (Ma et al., 2025), which uses Bézier curves to represent the full motion trajectory of objects, we employ the curve here to model the residual SE(3) deviation from the rigid FK prior. We parameterize the residual transformations as 9D vectors $\boldsymbol{\delta}=[\Delta\mathbf{x},\Delta\mathbf{r}]$, consisting of a 3D translation $\Delta\mathbf{x}$ and a continuous 6D rotation representation $\Delta\mathbf{r}$ (Zhou et al., 2019). The Bézier curve, defined by $K+1$ learnable control points $\{\mathbf{p}_{k}\}_{k=0}^{K}\subset\mathbb{R}^{9}$, outputs a 9D residual vector for any time $t$:
@@
where $B_{k}^{K}(t)$ are the Bernstein basis polynomials. We use a high-order curve ($K=19$) to capture complex motion patterns and scale the output by an influence factor $\omega$ to ensure it acts as a refinement. The resulting 9D vector is then converted to a valid SE(3) matrix $\mathbf{T}_{\text{Bézier}}(t)$ via Gram-Schmidt orthogonalization for the rotational part.
@@
To address the articulated structure of the robot, we introduce a static, per-joint offset $\mathbf{T}_{\text{embed}}^{(k)}$, represented by a learnable embedding $\mathbf{e}^{(k)}\in\mathbb{R}^{9}$ for each moving joint $k$. These time-invariant embeddings are optimized to correct for intrinsic structural discrepancies, such as link length errors or joint calibration drift. By integrating these local offsets into the kinematic chain alongside the global Bézier correction, our model specifically accounts for the hierarchical dependencies of robotic motion—a feature that distinguishes our approach from general, unstructured trajectory modeling.
@@
The final refined pose for each joint is obtained by integrating the nominal FK pose from MuJoCo with our two hierarchical corrections. The composition order is designed to handle errors at different levels of the kinematic chain: the global Bézier correction first compensates for temporal drifts or base misalignments, while the static embedding further refines the local joint-to-link transformation. For a specific joint $k$, the final world-space transformation is formulated as
$\mathbf{T}_{\text{final}}^{(k)}(t)=\mathbf{T}_{\text{FK}}^{(g)}(t)\circ\mathbf{T}_{\text{Bézier}}(t)\circ\mathbf{T}_{\text{embed}}^{(k)}$ (7)
where $\mathbf{T}_{\text{FK}}^{(g)}(t)$ denotes the global pose derived from the URDF model. This unified formulation ensures that the digital asset remains kinematically sound while precisely aligning with the visual observations in the video.
@@ 3.4 Optimization and Regularization
The primary supervision signal is a rendering loss comparing the rendered images with ground truth frames. Following standard practice (Kerbl et al., 2023), we use a combination of an L1 photometric loss and a D-SSIM term:
@@
To enforce the topological constraints of SGB module and ensure Gaussians remain well-aligned with their parent mesh faces, we adopt two regularization terms inspired by (Qian et al., 2024). We penalize the local position offset $\mu_{j}$ and scale $s_{j}$ if they exceed predefined thresholds, preventing primitives from detaching from the rigid structure or exhibiting visual jitter:
@@ 4 Experiments
Table 1: Quantitative comparison. Our model demonstrates superior performance in both static Novel-View Synthesis (left) and dynamic Novel-Pose Synthesis (right). Green indicates the best and yellow indicates the second-best performance.
@@ 4.1 Setup
Existing public robotics datasets often lack the precise alignment between visual observations and kinematic states required for high-fidelity asset creation. To facilitate research in this area, we present RoboArm4D, a benchmark designed for the generation and evaluation of dynamic digital assets. This dataset features several widely-used industrial arms, including the Franka Research 3, UR-5e, and ABB IRB 120. For each platform, we provide synchronized data packages comprising monocular video sequences, calibrated camera parameters, time-stamped joint trajectories, and the corresponding URDF files. This collection is intended to support the development of methods that reconcile rigid kinematic priors with actual physical motion. Further details regarding the data collection and package structure are available in the appendix.
@@
Our model, RoboArmGS, is implemented in PyTorch and trained for 600,000 iterations on a single NVIDIA H100 GPU using the Adam optimizer. For 3D Gaussian attributes, we follow the standard learning rate schedule from 3DGS (Kerbl et al., 2023), with position learning rates exponentially decayed to 1% of initial values. The BMR parameters are optimized with a learning rate of 0.0015, weight decay of 0.0001, and influence factor $\omega=0.1$. Gaussians are initialized by uniform sampling on the kinematic mesh surface, with binding-aware adaptive densification activated every 100 iterations from iteration 500 to 60,000; unstable Gaussian opacities are reset every 3,000 iterations to prune floaters. FK is computed via MuJoCo (Todorov et al., 2012) engine, and regularization hyperparameters are set as $\lambda_{\text{pos}}=0.01,\epsilon_{\text{pos}}=1.0$, $\lambda_{\text{scale}}=1.0,\epsilon_{\text{scale}}=0.6$, and $\lambda_{\text{vel}}=0.001$.
Figure 3: Qualitative comparison on novel-view and novel-pose synthesis. RoboArmGS achieves photorealistic rendering with precise geometric alignment across both tasks. While baseline methods suffer from severe blurring, structural distortion, or rendering artifacts under unseen configurations, our approach maintains high fidelity and kinematic accuracy, ensuring sharp and coherent results.
@@ 4.2 Main Results
We first evaluate static 3D reconstruction fidelity through the novel-view synthesis task. As reported in Tab. 1, RoboArmGS achieves state-of-the-art results across all metrics and significantly outperforms other methods. Although Robo-GS (Lou et al., 2025) is a robot-specific method, it still relies on the vanilla 3D Gaussian model for rendering, which performs unconstrained spatial optimization. In contrast, our SGB module explicitly anchors Gaussians to the mesh surface to incorporate the robot’s kinematic structure as a strong geometric prior. This structured approach prevents the formation of floating artifacts and maintains strict structural coherence, providing a more robust foundation for dynamic modeling. Qualitative comparisons in Fig. 3 further demonstrate that our method reconstructs fine-grained surface details with higher visual fidelity than these unconstrained baselines.
@@
We evaluate dynamic modeling capability through the novel-pose synthesis task, which involves rendering the robot in unseen joint configurations. As shown in Tab. 1, RoboArmGS substantially outperforms all baselines across all metrics. Generic 4D reconstruction methods (Wu et al., 2024; Yang et al., 2024) model temporal evolution using time indices rather than kinematic parameters. Consequently, they lack the ability to render from specified joint angles, a capability that is essential for digital assets applications. Even on held-out test frames, these generic deformation models fail to capture complex articulated motions and produce significant artifacts. To evaluate kinematic-driven approaches, we implement a 3DGS + FK baseline. This baseline binds 3D Gaussians to the mesh and drives their motion via nominal forward kinematics without any motion refinement. While Robo-GS (Lou et al., 2025) shares a similar philosophy, its reliance on manual part segmentation and intensive parameter tuning limits its scalability. Therefore, we utilize 3DGS + FK as the representative kinematic baseline. Although this approach maintains the robot’s rigid structure, it produces significant rendering errors, particularly in color and texture fidelity. These discrepancies arise because the nominal URDF poses deviate from the physical reality in the video, causing the rendered Gaussians to misalign with the actual image content. Our Bézier-based Motion Refiner addresses this by integrating the refinement of robotic motion directly into the 3D Gaussian optimization loop. By jointly optimizing the Bézier parameters alongside Gaussian attributes, our model reconciles nominal kinematics with actual visual dynamics through temporally smooth corrections. As demonstrated in Fig. 3, RoboArmGS produces sharp, structurally coherent images that align precisely with the ground truth, validating its accuracy as a controllable digital asset.
@@ 4.3 Ablation Study
[…]
    - w/o Adaptive Densification 20.414 0.803 0.451     - w/o Bézier Correction 29.384 0.956 0.052
@@
We validate the contribution of each proposed component through comprehensive ablation studies, as presented in Tab. 2. For the novel-view synthesis task, the removal of the SGB module leads to a severe degradation in reconstruction quality. Further ablating its sub-components, such as adaptive densification and binding mechanisms, confirms their necessity for achieving high-fidelity static representation. Regarding the novel-pose synthesis task, the BMR module plays a critical role in reconciling kinematic discrepancies. As illustrated in Fig. 4, the impact of BMR is highly motion-dependent. In regions with minimal movement, such as the robot base (green box), the differences between the full model and the ablated variant are negligible. Conversely, in regions experiencing large-scale articulation, such as the gripper (red box), the absence of BMR results in significant visual artifacts and misalignments. These results demonstrate that the learnable Bézier correction for global dynamics and the per-joint static offsets for local errors are complementary. This study empirically validates that all components of SGB and BMR are indispensable for the photorealistic and kinematically accurate performance of RoboArmGS.
@@
Figure 4: Ablation study visualization. Low-motion regions (green) show consistent results, whereas highly articulated components (red) exhibit artifacts without BMR, underscoring its necessity for complex motion alignment.
@@
We examine the sensitivity of the influence factor $\omega$, which modulates the learnable Bézier correction. The results for the novel-pose synthesis task are summarized in Tab. 3. Setting $\omega=0.0$ effectively disables the global motion refinement, which yields the lowest performance and confirms that nominal kinematics alone are insufficient for precise physical alignment. Even a minimal correction factor of $\omega=0.01$ leads to a substantial improvement across all metrics. The performance peaks at $\omega=0.1$, providing the optimal degree of refinement for aligning the digital asset with visual observations. However, further increasing $\omega$ to $1.0$ results in a slight performance decline, suggesting that an excessively large correction magnitude may introduce training instability or overfitting to specific motion trajectories. Consequently, we select $\omega=0.1$ as the standard setting to ensure a robust balance between motion accuracy and stability.
@@ 5 Conclusion
In this paper, we present RoboArmGS, a hybrid representation that synergizes Structured Gaussian Binding (SGB) with a Bézier-based Motion Refiner (BMR) to achieve both geometric consistency and precise motion modeling. By anchoring Gaussians to the robot’s kinematic structure and explicitly modeling physical motion residuals, our method effectively reconciles the discrepancy between nominal URDF models and real-world observations. To support future research, we contribute RoboArm4D, a specialized benchmark designed for the reconstruction and evaluation of dynamic robotic digital assets. Extensive experiments demonstrate that RoboArmGS achieves state-of-the-art performance in both novel-view and novel-pose synthesis, providing a robust and scalable solution for building photorealistic, kinematically accurate digital assets.
@@ B.2.1 Camera Pose Calibration
Accurately determining the pose of a static camera is crucial. To achieve this, we first captured a short calibration video (approx. 30 seconds) where the handheld camera was moved extensively around the static robotic arm and its environment. This multi-view sequence was processed using VGGT (Wang et al., 2025c) to generate a sparse 3D reconstruction of the scene and to precisely compute the camera poses for each frame of the calibration video. From this set of calibrated poses, we selected a single, fixed viewpoint for the subsequent motion capture. This process effectively pre-calibrates the static camera’s extrinsic parameters within the scene’s coordinate frame.
@@ B.2.3 Data Processing Pipeline
  • Camera Pose Estimation: We used VGGT (Wang et al., 2025c) to estimate the camera intrinsics and the per-frame extrinsic poses (camera-to-world transformation). Poses were optimized over the entire sequence to ensure global consistency.
  • Foreground Segmentation: To separate the robotic arm from the background, we employed the Segment Anything Model 2 (SAM2) (Ravi et al., 2024). We provided a few initial keyframe masks, and SAM2 (Ravi et al., 2024) automatically propagated the segmentation to the entire sequence, followed by minor manual refinement where necessary. This resulted in a pixel-perfect foreground mask for each frame.
@@ Appendix C Additional Experiments Results
In this section, we provide a detailed per-scene quantitative analysis of RoboArmGS on the RoboArm4D dataset, specifically focusing on the Universal Robots UR5e and ABB IRB 120 sequences. As shown in Tab. 4, our method achieves consistently superior performance across different robotic morphologies, significantly outperforming existing state-of-the-art (SOTA) methods and validating the generalizability of our proposed framework.
Novel-View Synthesis Novel-Pose Synthesis
[…]
2DGS (Huang et al., 2024a) 29.972 0.939 0.074 3DGS + FK 26.086 0.945 0.056
Robo-GS (Lou et al., 2025) 29.811 0.938 0.074 4DGS (Wu et al., 2024) 28.239 0.967 0.039
[…]
Table 4: Quantitative comparison on UR5e and ABB We evaluate our method against baselines on both Novel-View Synthesis (left) and Novel-Pose Synthesis (right). The top section shows results for the UR5e arm, and the bottom section for the ABB arm.